# Concurrent Fermionic Simulation Gate

### v1

@@ I Introduction
Quantum computing has made remarkable progress, achieving milestones such as quantum error correction, which enables practical computation despite challenges like decoherence [1]. Modern quantum processors now support around 100 qubits, and quantum supremacy demonstrations have shown limited speedups [2, 3]. As highlighted recently in Ref. [4], the next goal is achieving quantum utility by efficiently solving practical problems. This requires significant advancements in the challenges for 100-1M physical qubit processors, such as qubit coherence, gate accuracy and compactification [4, 5].
A quantum processor is considered universal if it can simulate any quantum circuit to arbitrary accuracy using a finite set of gates acting on a finite number of qubits. This requires single-qubit Hamiltonians to generate all $SU(2)$ gates and at least one two-qubit interaction to produce entangling gates [6, 7, 8, 9]. Discussions of universality often focus on unitary gates (exponentiated Hamiltonians) rather than physical Hamiltonians.
Recent technological advancements for controlling non-unitary errors, e.g. nonequilibrium quasiparticle tunneling [10, 11, 12, 13] and trap state noise[14, 15], as well as unitary errors, e.g. $ZZ$ stray couplings [16, 17, 18], have improved gate fidelity, architecture, and manufacturing, have enabled problem-specific algorithms to leverage quantum computing’s power [19, 3, 20, 21, 22, 23, 24, 25]. These developments have set the stage for high-fidelity computing with thousands of qubits. Two most applicable two-qubit gates are iswap (swapping $|01\rangle\leftrightarrow|10\rangle$) and cphase (applying a $\pi$ phase to $|11\rangle$). Often, a small cphase is a byproduct of iswap and vice versa. However, it has been shown recently that one can better control the cphase byproduct of iswap to define a new gate, namely ‘fermionic Simulation (fSim).’ This new gate can be separated into two gates in certain conditions.
The fSim gate is a versatile quantum computing tool, particularly useful for simulating phase transitions in strongly correlated fermionic systems [26, 27, 28]. Combining the functionality of two gates reduces resource demands and enhances flexibility for parameter-specific applications.
The fSim gate has been implemented in superconducting qubits using bichromatic modulation of frequency detuning and coupling strengths or hybrid methods with parametric drives [29, 30]. These approaches enable continuous transformations, swapping $|01\rangle$ and $|10\rangle$ while introducing a controllable phase for $|11\rangle$. The gate’s parameters $\theta$ and $\varphi$ allow state swaps and apply a phase factor $\exp(i\varphi)$, as represented by:
@@
Figure 1: Representation of 4 fermions $|\psi_{1},\psi_{2},\psi_{3},\psi_{4}\rangle$ on a column of 4 qubits $Q_{a},\ Q_{b},\ Q_{c}$, and $Q_{d}$. It takes six two-qubit concurrent fSim (cfSim) gates to transform the initial state into $|\psi_{4},\psi_{3},\psi_{2},\psi_{1}\rangle$.
@@
We propose an efficient theoretical modification to the fSim gate that integrates iswap and phase operations into a single hardware. This enables their simultaneous and controllable application, namely the “concurrent fSim (cfSim).” This approach compresses the gate time of consecutive iswap and cphase operations into one shorter gate with the same output. Our theoretical study demonstrates that this method simultaneously achieves the full range of iswap angles and conditional phases, inherently facilitating a fermion swap in a single operation. For its realization, we employ the example of bichromatic parametric drives at distinct frequencies in a transmon-coupler-transmon setup. The method uses two parametric drives: one resonating with the $|01\rangle\leftrightarrow|10\rangle$ transition to control the iswap angle, and the other facilitating cphase transitions via $|11\rangle\leftrightarrow|02\rangle$, $|11\rangle\leftrightarrow|20\rangle$, or both. Completing an oscillation cycle induces a detuning-dependent phase shift in the $|11\rangle$ state and reduces leakage at the same time. Moreover, we develop precise analytical strategies to accurately predict transition dynamics and resolve crosstalk between the drives, resolving synchronization issues.
II The Concept
One key application of quantum computing is simulating many-body models. A crucial aspect of state preparation for these systems involves constructing a adjacent fermionic swap network to ensure that many-body states satisfy the anticommutation relations of fermionic states.
Let $|\psi_{1}\cdots\psi_{i}\cdots\rangle$ represent the state of distinguishable and labeled fermions, where $\psi_{i}$ denotes the state of the $i$-th fermion. Swapping the states of the first and second fermions in $|\psi,\phi\rangle$ not only exchanges their states but also introduces a $\pi$-rotation phase, i.e., $e^{i\pi}|\phi,\psi\rangle$. More generally, the fermionic swap operator ${\hat{\textup{S}}}_{ij}$ exchanges the states of the $i$-th and $j$-th fermions with a phase: ${\hat{\textup{S}}}_{ij}|\psi_{1}\cdots\psi_{i}\cdots\psi_{j}\cdots\rangle=e^{i\pi}|\psi_{1}\cdots\psi_{j}\cdots\psi_{i}\cdots\rangle$. Applying ${\hat{\textup{S}}}_{ij}$ twice restores the original state, i.e., ${\hat{\textup{S}}}_{ij}{\hat{\textup{S}}}_{ji}=1$. Additionally, as a unitary operator, ${\hat{\textup{S}}}_{ij}$ satisfies $({\hat{\textup{S}}}_{ij})^{\dagger}={\hat{\textup{S}}}_{ji}$.
For tensor networks in classical simulations of fermionic systems [34], the fermionic swap operator can be expressed independently of any specific qubit mapping of fermions. It is defined using the fermionic operators $\hat{f}^{\dagger}_{i}$ (creation) and $\hat{f}_{i}$ (annihilation), which act on the vacuum state $|{\rm vac}\rangle_{i}$ at site $i$ and satisfy the anticommutation relation $\{\hat{f}_{i},\hat{f}^{\dagger}_{j}\}=\delta_{ij}$. The swap operator satisfies ${\hat{\textup{S}}}_{ij}\hat{f}^{\dagger}_{i}{\hat{\textup{S}}}^{\dagger}_{ij}=\hat{f}^{\dagger}_{j}$ and ${\hat{\textup{S}}}_{ij}\hat{f}_{i}{\hat{\textup{S}}}^{\dagger}_{ij}=\hat{f}_{j}$. This leads to the explicit form ${\hat{\textup{S}}}_{ij}=\hat{f}^{\dagger}_{i}(\hat{f}_{j}-\hat{f}_{i})+\hat{f}^{\dagger}_{j}(\hat{f}_{i}-\hat{f}_{j})+1$, preserving the anticommutation properties of fermions.
Qubit mapping is a useful transformation representing multi-particle fermionic states using qubits, such as the Jordan-Wigner transformation [35]. This mapping relies on the ordering of operators and qubits.
For instance, fermion states $|\psi_{i}\rangle$ can be encoded as qubit states $|0\rangle_{i}$ and $|1\rangle_{i}$ at site $i$ with the mapping:
${\rm fermion}\ |{\rm vac}\rangle_{i}$$\leftrightarrow$ ${\rm qubit}\ |0\rangle_{i}$
${\rm fermion}\ |\psi_{i}\rangle$$\leftrightarrow$ ${\rm qubit}\ |1\rangle_{i}$ (2)
with $|\psi_{i}\rangle$ being $\hat{f}^{\dagger}_{i}|{\rm vac}\rangle_{i}$.
Consider the general fermionic Hamiltonian: $H=\sum_{ij}A_{ij}\hat{f}^{\dagger}_{i}\hat{f}_{j}+\sum_{i}B_{i}\hat{n}_{i}+\sum_{i\neq j}C_{ij}\hat{n}_{i}\hat{n}_{j}$, with $\hat{n}_{i}=\hat{f}^{\dagger}_{i}\hat{f}_{i}$ as the number operator. This Hamiltonian appears in models like the Hubbard model and quantum chemistry, with non-local hopping terms $A_{ij}$ and four-body interaction terms $C_{ij}$. Importantly, $H$ remains unchanged under ${\hat{\textup{S}}}_{ij}$, making ${\hat{\textup{S}}}_{ij}$ a symmetry operator. Time evolution under $H$ can thus incorporate the fermionic swap.
The fermionic swap exchanges $\psi_{i}$ and $\psi_{j}$ while maintaining proper antisymmetrization. This is significant for qubit representations of fermionic operators under the Jordan-Wigner transformation, where the ordering follows a canonical sequence. This mapping fractionates ${\hat{\textup{S}}}_{ij}$ into multiple swap gates between adjacent qubits [34].
Under the Jordan-Wigner transformation, the corresponding qubit Hamiltonian follows canonical ordering, requiring gates between neighbouring qubits. This composite two-qubit gate defines the fermionic simulation (fSim) gate. Explicitly:
$\widehat{\textup{fSim}}={\rm\mathbb{P}}(e^{-iC\hat{n}_{i}\hat{n}_{i+1}t}\ e^{-iA(\hat{f}^{\dagger}_{i}\hat{f}_{i+1}+h.c.)t}{\hat{\textup{S}}}_{i,i+1})$ (3)
$A\equiv A_{i,i+1}$ and $C\equiv C_{i,i+1}$ are the corresponding interactions between a pair of adjacent qubits. $\mathbb{P}$ denotes the projection onto the qubit subspace defined in Eq.(2), Eq.(3) simplifies to the $4\times 4$ computational subspace:
$\widehat{\textup{fSim}}=\begin{pmatrix}1&0&0&0\\ 0&-i\sin(At)&\cos(At)&0\\ 0&\cos(At)&-i\sin(At)&0\\ 0&0&0&-ie^{-iCt}\end{pmatrix}$ (4)
Comparing Eq.(4) and Eq.(1) shows that controlling the parameters $A$, $C$, and the gate time $t$ are crucial for tuning the gate angles $\varphi$ and $\theta$, They turn $\theta=At-\frac{\pi}{2}$ and $\varphi=\pi/2-Ct$, up to single qubit rotations.
Fig.(1) illustrates an example with four fermion states $|\psi_{1}\psi_{2}\psi_{3}\psi_{4}\rangle$ represented by the vertical column of four qubits $Q_{a},\ Q_{b},\ Q_{c},$ and $Q_{d}$ shown by circle of distinct colors. The typical fermionic state $\psi_{i}$ is denoted by $|i\rangle$ inside each circle. After four layers of cfSim operations, i.e. vertical dashed line, the output fermion state on the qubits will correspond to the swapped state $|\psi_{4}\psi_{3}\psi_{2}\psi_{1}\rangle$.
III Concurrent fermionic simulation gate
@@
In superconducting circuits consisting of two transmons coupled via a frequency-tunable coupler, a continuous two-qubit fSim gate has recently been realized. Indivibichromatic iswap and cphase gates can be implemented by adjusting the qubit-qubit coupling strength and frequency detuning, with one gate activated immediately after the other is deactivated [29]. Notably, both gates can also be active simultaneously, enabling the complete and concurrent implementation of the fermionic simulation gate set in a single operation, referred to as concurrent fermionic simulation (cfSim). Below, we present the theory for realizing such a gate using two parametric drives in a controllable manner.
III.1 Setup: bichromatic parametric driving (BPD)
@@
Motivated by the idea of realizing fSim gate by a parametric drive [36, 37, 38], in this paper, we consider the simultaneous application of two gates achieved by employing Bichromatic Parametric Driving (BPD). In this approach, we apply two parametric drives at distinct frequencies, each on one qubit. The schematic of the setup is depicted in Fig.(3). The first drive resonates the state $|11\rangle$ with either $|02\rangle$ or $|20\rangle$. The second drive facilitates a swap between the single-qubit excited states $|10\rangle$ and $|01\rangle$ by bringing them into resonance, as illustrated in Fig.(3). The state $|00\rangle$ remains uncoupled from other states because it is far detuned from the relevant transitions, thereby staying unaffected by the applied drives.
@@
The qubit dynamics of an iswap /cphase gate set can be efficiently modeled by two interacting qutrits with time-depending frequency detuning $\Delta(t)=\omega_{1}(t)-\omega_{2}(t)$ and time-depending coupling strength $g(t)$. The Hamiltonian of such a system in the basis of $|01\rangle$, $|10\rangle$, $|11\rangle$, $|20\rangle$, $|02\rangle$ from left (top) to right (bottom) is given by:
$H_{0}=\begin{pmatrix}0&g&0&0&0\\ g&\Delta&0&0&0\\ 0&0&\Delta&\sqrt{2}g&\sqrt{2}g\\ 0&0&\sqrt{2}g&2\Delta+\delta&0\\ 0&0&\sqrt{2}g&0&\delta\end{pmatrix}$ (5)
@@
Figure 2: Setup of the system. Two transmons are coupled via a common harmonic resonator. The direct transmon-transmon coupling is ignored. Two parametric drives are applied to the two transmons separately. There are five control parameters of the drives: two drive amplitudes $\Omega_{1}$ and $\Omega_{2}$, two drive frequencies $\nu_{1}$ and $\nu_{2}$ and one relative phase $\phi_{1}^{d}-\phi_{2}^{d}$. In our analysis, the relative phase is fixed to be zero. The two simultaneous drives induce a concurrent fSim gate between the two qubits, exchanging their excitations and adding a conditional phase $\varphi$.
Our proposed gate operates by modulating the frequencies $\omega_{1}$ and $\omega_{2}$ in an oscillatory manner. Consider the setup depicted in Fig.(3), where both transmons $Q_{1}$ and $Q_{2}$, coupled via a resonator, are parametrically driven by $h_{i}(t)\equiv\Omega_{i}\sin(\nu_{i}t+\phi_{i}^{d})\hat{a}^{\dagger}_{i}\hat{a}_{i}$ for $i=1,2$. Here, $\phi_{i}^{d}$ is the driving phase, $\nu_{i}$ the modulation frequency, and $\Omega_{i}$ the modulation amplitude.
Since a parametric drive modifies the qubit frequency, it effectively acts as a time-dependent number operator $\hat{a}^{\dagger}_{i}\hat{a}_{i}$, where $\hat{a}^{\dagger}_{i}$ ($\hat{a}_{i}$) represents the creation (annihilation) operator for the qutrit $i$. The total Hamiltonian within the $5\times 5$ qutrit Hilbert space of interest is expressed as:
$H=H_{0}+\begin{pmatrix}h_{2}(t)&0&0&0&0\\ 0&h_{1}(t)&0&0&0\\ 0&0&\sum_{i}h_{i}(t)&0&0\\ 0&0&0&2h_{1}(t)&0\\ 0&0&0&0&2h_{2}(t)\end{pmatrix}$ (6)
Figure 3: Schematic of BPD for cfSim gates. Relevant levels are shown. Two parametric drives target two transitions separately. When on resonance, drive frequency $\nu_{1,2}$ should match level detuning. Drive amplitude $\Omega_{1,2}$ are not required to match detunings. For cfSim gates, we choose to drive 1 to resonantly drive $|01\rangle\leftrightarrow|10\rangle$ transition, therefore, $\nu_{1}=\omega_{2}-\omega_{1}$. The second drive is chosen to be near resonant with $|11\rangle\leftrightarrow|20\rangle$ or $|11\rangle\leftrightarrow|02\rangle$ transition. iswap and cphase transitions are happening simultaneously. Both iswap angle $\theta$ and conditional phase $\varphi$ can be controlled.
[…]
@@
in which $J_{n}$ is the Bessel function of the first kind, and the summation takes place over integers $n\in\mathbb{Z}$.
Without loss of generality, let us assume that the frequency of Q1 is greater than that of Q2, i.e. $\omega_{1}>\omega_{2}$. For this setup, we can parametrically drive the frequency of Q1 by supplying the required energy to achieve the transition $|01\rangle\leftrightarrow|10\rangle$, i.e. $\nu_{1}=\Delta$. The second qubit Q2 is parametrically driven to come very close to the resonance between $|11\rangle\leftrightarrow|02\rangle$ transition, i.e. $\nu_{2}=\Delta+\delta+\varepsilon$ with small $|\varepsilon|$, i.e. $|\varepsilon|\ll\min\{|\Delta+\delta|,|\Delta-\delta|\}$. By this assumption, only the slow-varying terms proportional to $\exp(i\varepsilon t)$ contribute to the state evolution in the long run $t\gg 1/(\Delta\pm\delta)$. The Hamiltonian (7) contains interaction within and outside the Hilbert space’s computational subset. Within the computational subset there is exchange interaction $|10\rangle\langle 01|$ with the strength $gJ_{1}(\Omega_{1}/\Delta)J_{0}(\Omega_{2}/(\Delta+\delta+\varepsilon))$. Outside of the computational subset the energy level $|11\rangle$ interacts with $|02\rangle$ and $|20\rangle$ by the following coupling strength $\sqrt{2}gJ_{0}(\Omega_{1}/\Delta)J_{1}(\Omega_{2}/(\Delta+\delta+\varepsilon))e^{i\varepsilon t}$ and $\sqrt{2}gJ_{-2}(\Omega_{1}/\Delta)J_{1}(\Omega_{2}/(\Delta+\delta+\varepsilon))e^{i\varepsilon t}$, respectively.
From here on, this analysis can be performed numerically to simulate the behavior of the two qubits under the two parametric drives, however before that, we approach the problem perturbatively.
III.2 Weak BPD
It is helpful that we first consider some restrictions on the parameter domains to make the problem eligible for a valid solution in the perturbative orders.
One of the aspects of perturbatively validity of our result is that the qubits are weakly driven, therefore we consider that the drive amplitude is weak, i.e. ${\Omega_{i}}<\nu_{i}$ for $i=1,2$. Within this limit which makes the arguments of the Bessel function, small, one can show that $J_{-2}J_{1}\ll J_{1}J_{0}$; therefore, the coupling strength of $|11\rangle\langle 20|$ transition turns out to be negligible compared to the other two interaction strengths and can be approximately ignored. Therefore in the interaction picture after rotating-wave-approximation, the BPD Hamiltonian can be simplified as
@@ III.2 Weak BPD
The residual time dependence in the coupling between $|11\rangle$ and $|02\rangle$ can be treated by applying another rotating frame transformation defined as $U_{R}=e^{-i\varepsilon t|02\rangle\langle 02|}\otimes I$ which rotates $|02\rangle$ state and leaves other levels unchanged. This helps to transform the interaction to the frame rotating with the frequency of level detuning between $|11\rangle$ and $|02\rangle$. The resulting Hamiltonian turns out to be block-diagonal with the following two diagonal blocks, that resemble separate iswap gates in the subspace spanned by $|10\rangle,|01\rangle$ and cphase gate in the subspace spanned by $|11\rangle,|02\rangle$, as follows:
$H_{\rm{iswap}}=\begin{pmatrix}0&gJ_{1}(x_{1})J_{0}(x_{2})\\ gJ_{1}(x_{1})J_{0}(x_{2})&0\\ \end{pmatrix},$ (9)
$H_{\rm{cphase}}=\begin{pmatrix}0&\sqrt{2}gJ_{0}(x_{1})J_{1}(x_{2})\\ \sqrt{2}gJ_{0}(x_{1})J_{1}(x_{2})&-\varepsilon\\ \end{pmatrix}$ (10)
Note that in our numerical analysis we carry on the analysis in the presence of all possible transitions and therefore we consider the weak driving limit of this section only to make analysis simpler for analytical verification. Performing the analytical study in the relatively intermediate regime of driving amplitude requires to reconsider the coupling $|11\rangle\langle 20|$ that needs to deal with $3\times 3$ matrix size everywhere.

### v2

@@ I Introduction
Quantum computing has made remarkable progress, achieving milestones such as quantum error correction, which enables practical computation despite challenges like decoherence [1, 2, 3]. Modern quantum processors now support around 100 qubits, and quantum supremacy demonstrations have shown limited speedups [4, 5]. As highlighted recently in Ref. [6], the next goal is achieving quantum utility by efficiently solving practical problems. This requires significant advancements in the challenges for 100-1M physical qubit processors, such as qubit coherence, gate accuracy and compactification [6, 7].
A quantum processor is considered universal if it can simulate any quantum circuit to arbitrary accuracy using a finite set of gates acting on a finite number of qubits. This requires single-qubit Hamiltonians to generate all $SU(2)$ gates and at least one two-qubit interaction to produce entangling gates [8]. Discussions of universality often focus on unitary gates rather than physical Hamiltonians.
Recent technological advancements for controlling non-unitary errors, e.g. nonequilibrium quasiparticle tunneling [9, 10, 11, 12] and trap state noise[13, 14], as well as unitary errors, e.g. $ZZ$ stray couplings [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], have improved gate fidelity, architecture, and manufacturing, have enabled problem-specific algorithms to leverage quantum computing’s power [27, 5, 28, 29, 30, 31, 32]. These developments have set the stage for high-fidelity computing with thousands of qubits. Two of the most commonly used two-qubit gates are iSWAP (swapping $|01\rangle\leftrightarrow|10\rangle$) and CPHASE (applying a $\pi$ phase to $|11\rangle$). Often, a small CPHASE is a byproduct of iSWAP and vice versa. However, it has been shown recently that one can better control the CPHASE byproduct of iSWAP to define a new gate, namely ‘fermionic Simulation (fSim).’ This new gate can be separated into two gates in certain conditions.
The fSim gate is a versatile quantum computing tool, particularly useful for simulating phase transitions in strongly correlated fermionic systems [33, 34, 35, 36]. Combining the functionality of two gates reduces resource demands and enhances flexibility for parameter-specific applications.
The fSim gate has been implemented in superconducting qubits using bichromatic modulation of frequency detuning and coupling strengths or a DC pulse with a parametric drive [37, 38]. These approaches enable continuous transformations, swapping $|01\rangle$ and $|10\rangle$ while introducing a controllable phase for $|11\rangle$. The gate’s parameters $\theta$ and $\varphi$ allow state swaps and apply a phase factor $\exp(i\varphi)$, as represented by:
@@
As an example, consider the general fermionic Hamiltonian: $H=\sum_{ij}A_{ij}\hat{f}^{\dagger}_{i}\hat{f}_{j}+\sum_{i}B_{i}\hat{n}_{i}+\sum_{i\neq j}C_{ij}\hat{n}_{i}\hat{n}_{j}$, with $\hat{n}_{i}=\hat{f}^{\dagger}_{i}\hat{f}_{i}$ as the number operator. This Hamiltonian appears in models like the Hubbard model and quantum chemistry, with hopping terms $A_{ij}$, on-site energy $B_{i}$, and electron-electron interaction $C_{ij}$. In quantum simulation, the on-site energy is mapped to single-qubit gates. The hopping terms correspond to iSWAP operations. The electron-electron interaction together with fermionic anti-commutator can be mapped to the conditional phase $\varphi$. Therefore, being able to implement iSWAP and CPHASE operations simultaneously makes simulating such kind of Hamiltonians more efficient.
Although the fSim concept offers several advantages for quantum algorithm design, some challenges hampered previous fSim implementations. In experiments, calibrating the gate in a continuous 2D parameter space is very time-consuming. From the theory side, there is no well-established theory for driving iSWAP transition and CPHASE transition simultaneously. One of the reasons is that for systems where two-qubit gates are activated via level detunings and static couplings, detunings and couplings in the one-excitation subspace (between $|01\rangle$ and $|10\rangle$) and two-excitation subspace (among $|11\rangle$, $|02\rangle$ and $|20\rangle$) can not be controlled independently, which makes the iSWAP transition and CPHASE transition correlated[37]. For systems where two-qubit gates are driven via fast oscillating pulses ( microwave drives or parametric flux drives), the lack of understanding of drive-drive crosstalk makes it difficult to control simultaneous transitions.
@@
To address those issues in theory, we propose a theoretical modification to the fSim gate that integrates iSWAP and CPHASE operations into a single hardware. This enables their simultaneous and controllable application, namely the “concurrent fSim (cfSim).” This approach compresses the gate time of consecutive iSWAP and CPHASE operations into one shorter gate with the same output. Our theoretical study demonstrates that this method simultaneously achieves the full range of iSWAP angles and conditional phases, inherently facilitating a fermion swap in a single operation. We note that there have been several previous works in recent years on concurrent driving of gates to implement complicated gates[42, 43, 44, 45]. In fact, in a recent experiment, it is suggested to run iSWAP and CZ simultaneously through a bichromatic parametric drive on the coupler between two qubits[46]. In this paper, we give a detailed analysis and simulations of how concurrent fSim gates can be implemented. We employ the example of bichromatic parametric drives at distinct frequencies in a transmon-coupler-transmon setup. The method uses two parametric drives: one resonating with the $|01\rangle\leftrightarrow|10\rangle$ transition to control the iSWAP angle, and the other facilitating CPHASE transitions via $|11\rangle\leftrightarrow|02\rangle$, $|11\rangle\leftrightarrow|20\rangle$, or both. Completing an oscillation cycle induces a detuning-dependent phase shift in the $|11\rangle$ state and reduces leakage at the same time. Moreover, we develop precise analytical strategies to accurately predict transition dynamics and resolve crosstalk between the drives, resolving synchronization issues.
The remainder of the paper is organized as follows. In Sec. II, we introduce the setup of bichromatic parametric drive and study two models based on it to understand the dynamics of two-tone drives. We first present analytical derivations of a two-qutrit toy model and then extend it to the more complete coupled Kerr nonlinear oscillator model. Section III deals with an important aspect of superconducting quantum gates: leakage. We show how to minimize leakage when two drives are present at the same time. In Section IV, we combine all analyses above to derive analytical expressions of iSWAP angle $\theta$ and conditional phase $\varphi$ in cfSim gates. Numerical simulation and optimization of cfSim are performed. Sec. V presents an important case of cfSim gates: ZZ-free iSWAP gate. In Sec.VI, we discuss what limits the cfSim gate fidelity and give some outlook of our proposed gate scheme. In Section VII, we summarize this paper.
II The Model
@@
In superconducting circuits consisting of two transmons coupled via a frequency-tunable coupler, a continuous two-qubit fSim gate has recently been realized. iSWAP and CPHASE gates can be implemented by adjusting the qubit-qubit coupling strength and frequency detuning, with one gate activated immediately after the other is deactivated [37]. Notably, both gates can also be active simultaneously, enabling the complete and concurrent implementation of the fermionic simulation gate set in a single operation, referred to as concurrent fermionic simulation (cfSim). Below, we present the theory for realizing such a gate using two parametric drives in a controllable manner.
II.1 Setup: bichromatic parametric driving (BPD)
Motivated by the idea of realizing fSim gate by a parametric drive [47, 48, 49], in this paper, we consider the simultaneous application of two gates achieved by employing Bichromatic Parametric Driving (BPD). In this approach, we apply two parametric drives at distinct frequencies, each on one qubit. Our proposed gate operates by modulating the frequencies $\omega_{1}$ and $\omega_{2}$ in an oscillatory manner. Consider the setup depicted in Fig.(1), where both transmons $Q_{1}$ and $Q_{2}$, coupled via a resonator, are parametrically driven via modulating external phases $\varphi_{e1}$ and $\varphi_{e2}$.
Figure 1: Setup of the system. Two transmons are coupled via a common coupler. The direct transmon-transmon coupling is ignored. Two parametric drives are applied to the two transmons separately. The two simultaneous drives induce a concurrent fSim gate between the two qubits, exchanging their excitations and adding a conditional phase.
The full circuit Hamiltonian reads[50]:
$H=H_{1}+H_{2}+H_{c}+H_{int}$ (2)
where the transmon Hamiltonians, coupler Hamiltonian, and the coupling Hamiltonian are expressed in the canonical phase-charge basis as:
$\begin{split}H_{j}&=4E_{Cj}\mathbf{\hat{n}}_{j}^{2}-E_{Jj}\cos(\bm{\hat{\varphi}}_{j}+\varphi_{ej}(t))\\ H_{int}&=\Sigma_{jk}4E_{Cjk}\bm{\hat{n}}_{j}\bm{\hat{n}}_{k}\end{split}$ (3)
with $j,k=1,2\text{ and }c$.
The analytical analysis and numerical simulation of the full circuit Hamiltonian is quite complicated. To focus on the essential points of our gate scheme, we instead study two simplified models. We first introduce a toy model consisting of two interacting qutrits to understand the basics of the gate dynamics under BPD. We then turn to a more complete model including transmon higher levels and coupler degrees of freedom, which is more accurate yet can still be studied analytically. Exact numerical simulation of the full circuit model can be performed with the help of methods such as time-dependent Schrieffer-Wolff transformation or Floquet theory[50].
II.2 A Toy Model: Two Qutrits
@@
The schematic of the proposed gate scheme is depicted in Fig. (2). The first drive resonates the state $|11\rangle$ with either $|02\rangle$ or $|20\rangle$. The second drive facilitates a swap between the single-qubit excited states $|10\rangle$ and $|01\rangle$ by bringing them into resonance, as illustrated in Fig.(2). The state $|00\rangle$ remains uncoupled from other states because it is far detuned from the relevant transitions, thereby staying unaffected by the applied drives.
@@
The qubit dynamics of an iSWAP /CPHASE gate set can be efficiently modeled by two interacting qutrits with time-dependent frequency detuning $\omega_{1}(t)-\omega_{2}(t)=\Delta+h_{1}(t)-h_{2}(t)$ and fixed coupling strength $g$. The static Hamiltonian of such a system in the basis of $|01\rangle$, $|10\rangle$, $|11\rangle$, $|02\rangle$, $|20\rangle$ from left (top) to right (bottom) is given by:
$H_{0}=\begin{pmatrix}0&g&0&0&0\\ g&\Delta&0&0&0\\ 0&0&\Delta&\sqrt{2}g&\sqrt{2}g\\ 0&0&\sqrt{2}g&\delta&0\\ 0&0&\sqrt{2}g&0&2\Delta+\delta\end{pmatrix}$ (4)
@@
Since a parametric drive modifies the qubit frequency, it effectively acts as a time-dependent number operator $\hat{a}^{\dagger}_{i}\hat{a}_{i}$, where $\hat{a}^{\dagger}_{i}$ ($\hat{a}_{i}$) represents the creation (annihilation) operator for the qutrit $i$. For simplicity, we assume anharmonicities are not affected by drives. The total Hamiltonian within the $5\times 5$ qutrit Hilbert space of interest is expressed as[37, 47, 48]:
$H=H_{0}+\begin{pmatrix}h_{2}(t)&0&0&0&0\\ 0&h_{1}(t)&0&0&0\\ 0&0&\sum_{i}h_{i}(t)&0&0\\ 0&0&0&2h_{2}(t)&0\\ 0&0&0&0&2h_{1}(t)\end{pmatrix}$ (5)
Figure 2: Schematic of BPD for cfSim gates. Relevant levels are shown. Two parametric drives target two transitions separately. When on resonance, drive frequency $\nu_{1,2}$ should match level detuning. Drive amplitudes $\Omega_{1,2}$ are not required to match detunings. For cfSim gates, we choose to drive 1 to resonantly drive $|01\rangle\leftrightarrow|10\rangle$ transition, therefore, $\nu_{1}=\omega_{2}-\omega_{1}$. The second drive is chosen to be near resonant with $|11\rangle\leftrightarrow|20\rangle$ or $|11\rangle\leftrightarrow|02\rangle$ transition. iSWAP and CPHASE transitions are happening simultaneously. Both iSWAP angle $\theta$ and conditional phase $\varphi$ can be controlled.
Let’s assume the drives $h_{1}$ and $h_{2}$ take the form: $h_{i}=\Omega_{i}\sin(\nu_{i}t)$, where drive amplitudes and frequencies are time-independent. We can then move to a rotating frame where all diagonal terms are zero, see Appendix (A). In this rotating frame, the Hamiltonian is given by:
@@
where $\Delta_{0}$ is the static detuning between $|01\rangle$ and $|10\rangle$, $J_{n}$ is the Bessel function of the first kind, and the summation takes place over integers $n\in\mathbb{Z}$.
In Appendix (B), we show that, under certain perturbative conditions, the Hamiltonian can be further simplified:
@@
with $x_{k}=\frac{\nu_{k}}{\Omega_{k}}$.
The residual time dependence in the coupling between $|11\rangle$ and $|02\rangle$ can be treated by applying another rotating frame transformation defined as $U_{R}=e^{-i\varepsilon t|02\rangle\langle 02|}\otimes I$, which rotates $|02\rangle$ state and leaves other levels unchanged. This helps to transform the interaction to the frame rotating with the frequency of level detuning between $|11\rangle$ and $|02\rangle$. The resulting Hamiltonian turns out to be block-diagonal with the following two diagonal blocks, that resemble separate iSWAP gates in the subspace spanned by $|10\rangle,|01\rangle$ and CPHASE gate in the subspace spanned by $|11\rangle,|02\rangle$, as follows:
$H_{10,01}$ $=\begin{pmatrix}0&gJ_{1}(x_{1})J_{0}(x_{2})\\ gJ_{1}(x_{1})J_{0}(x_{2})&0\\ \end{pmatrix}$ (8)
$H_{11,02}$ $=\begin{pmatrix}0&\sqrt{2}gJ_{0}(x_{1})J_{1}(x_{2})\\ \sqrt{2}gJ_{0}(x_{1})J_{1}(x_{2})&-\varepsilon\\ \end{pmatrix}$ (9)
Note that in our numerical analysis, we carry on the analysis in the presence of all possible transitions, and therefore, we consider the weak driving limit of this section only to make analysis simpler for analytical verification. Performing the analytical study in the relatively intermediate regime of driving amplitude requires to reconsider the coupling $|11\rangle\langle 20|$ that needs to deal with $3\times 3$ matrix size everywhere.